OpenStudy (anonymous):

pam has 30 plant photots, 100 cat photos, and 70 dog photos. in a cicle graph, which angle measure would be used for the plant section?

OpenStudy (anonymous):

30 + 100 + 70 = 200 ' ' /2 15 + 50 + 35 = 100 360/15 = 54 degrees

OpenStudy (anonymous):

thanks